OpenStudy (anonymous):

Someone help with an algebra 2 problem!!!

OpenStudy (anonymous):

5f-(3f+4)=20 The parantheses stand for absolute value

OpenStudy (zale101):

Your question was posted on the wrong category. Go to http://openstudy.com/study#/groups/Mathematics and post ur question there

OpenStudy (anonymous):

-15f+-20=20 -15f=20+20 -15f=40 f=40*-15

OpenStudy (anonymous):

i'm not so sure